OpenStudy (anonymous):

What value of x in the interval 0

OpenStudy (anonymous):

start with \[\tan(x)=-\frac{1}{\sqrt3}\]

OpenStudy (anonymous):

then see on the unit circle where that is the value for tangent actually you will see the point on the unit circle where \(\cos(x)=\frac{\sqrt3}{2}\) and \(\sin(x)=\frac{1}{2}\)

OpenStudy (anonymous):

if you do not have a unit circle handy, look at the last page of the attached cheat sheet

OpenStudy (anonymous):

thank you!